OpenStudy (anonymous):

Input the equation of the given line in standard form. The line with m = 2/3 and passing through (1, 1)

OpenStudy (anonymous):

y-1=2/3(x-1) 2/3x-y=2/3-1=-1/3 so: 2x-3y=-1